OpenStudy (anonymous):

What is the product in lowest terms? 3/4 x 5/9 A. 5/36 B. 5/13 C. 5/12 D. 8/13

OpenStudy (anonymous):

3x5=? 4x9=?

OpenStudy (anonymous):

15 and 36

OpenStudy (anonymous):

so what divides into both 15 and 36?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

well 15 is an odd number, so 2 can't divide into 15 without having a remainder

OpenStudy (anonymous):

1

OpenStudy (anonymous):

okay, write out your 3 times table (up to 12x3), and tell me if both 15 and 36 are in it?

OpenStudy (anonymous):

they are not both in it

OpenStudy (anonymous):

write your 3x table out on here for me, so i can see where you are doing wrong :)

OpenStudy (anonymous):

3 x 1=3 3 x 2 =6 3 x 3=9 3 x 4=12 3 x 5=15 3 x 6= 18 3 x 7= 21 3 x 8 =24 3 x 9=27 3 x 10= 30 3 x 11 =33 3 x 12= 36

OpenStudy (anonymous):

i see where i went wrong

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

so it s 3

OpenStudy (anonymous):

so to put a fraction in its lowest terms, the goal is to find a common number that divides them both, so we have: \[\frac{ 3 }{ 4 }*\frac{ 5 }{ 9 }=\frac{ 15 }{ 36 }\] so if we divide the top and bottom numbers by '3' - becuase we know that they both divide, what do we get? (yes it is 3) :)

OpenStudy (anonymous):

so we have \[\frac{ 15 \div 3 }{ 36\div3 }= \frac{ ? }{ ? }\]

OpenStudy (anonymous):

I dun even know the answer, but I'd be happy for someone to mark me as Best Response for doing Absolutely Nothing!!

OpenStudy (anonymous):

oh wait, this question is easy

OpenStudy (anonymous):

5 and 4

OpenStudy (anonymous):

top one is right, 15/3=5 but 36/3 does not equal 4...look at the table you wrote about

OpenStudy (anonymous):

36 is bigger than 15 so the answer has to be bigger than 5

OpenStudy (anonymous):

12

OpenStudy (anonymous):

yes!! so if we have 15/3=5 on the top and 36/3=12 on the bottom what fraction do we have?

OpenStudy (anonymous):

5/12

OpenStudy (anonymous):

yes!! that is correct!

OpenStudy (anonymous):

thank you